the square root of (-3x-14) -x=4
Answer by mukhopadhyay(490)

Sqrt(-3x-14) - x = 4
=> Sqrt(-3x-14) = x+4
=> (-3x-4) = x^2 + 8x + 16 (squaring both sides)
=> x^2 +11x + 20 = 0
=> x = [-11 + sqrt(121-80)]/2 or x = [-11 - sqrt(121-80)]/2
=> x = [-11 + sqrt(41)]/2 or x = [-11 - sqrt(41)]/2